OpenStudy (anonymous):

An equation which represents the average weight in pounds of an adult male of a given height in inches is w = 4.5h - 142. Predict the weight in pounds of an average male with a height of 6' 3" (75 inches). Round to the nearest pound. A. 168 B. 182 C. 196 D. 212

Directrix (directrix):

Take the equation: w = 4.5h - 142 and replace h by 75 inches to get an estimate of the man's weight. w = 4.5h - 142 w = 4.5 * 75 - 142 w = ? After that, there is the rounding. @cassiesilva0628

Directrix (directrix):

In the Google search box, paste this: 4.5 * 75 - 142 = The Google calculator will pop up with the answer.

OpenStudy (anonymous):

C ? @Directrix

Directrix (directrix):

w = 195.5 which rounds up to 196 I agree with option C.

OpenStudy (anonymous):

okay thank you so much

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

do you have time for some other questions? @Directrix

Directrix (directrix):

Yes. I see that you have some good helpers already.